# TLSQL: Table Learning Structured Query Language

### v1

@@ Abstract.
Table learning, which lies at the intersection of machine learning and modern database systems, has recently attracted growing attention. However, existing frameworks typically require explicit data export and extensive feature engineering, creating a high barrier for database practitioners. We present TLSQL (Table Learning Structured Query Language), a system that enables table learning directly over relational databases via SQL-like declarative specifications. TLSQL is implemented as a lightweight Python library that translates these specifications into standard SQL queries and structured learning task descriptions. The generated SQL queries are executed natively by the database engine, while the task descriptions are consumed by downstream table learning frameworks. This design allows users to focus on modeling and analysis rather than low-level data preparation and pipeline orchestration. Experiments on real-world datasets demonstrate that TLSQL effectively lowers the barrier to integrating machine learning into database-centric workflows. Our code is available at https://github.com/rllm-project/tlsql/.
@@ 1. Introduction
In summary, TLSQL establishes a practical bridge between modern databases and machine learning by introducing a SQL-like declarative paradigm for table learning task specification. By lowering the barrier to entry, TLSQL enables database users to seamlessly integrate machine learning into existing SQL-centric workflows. We believe that TLSQL offers both a novel design perspective and a practical system prototype, paving the way for future database-centric learning systems.
@@ 3.3. Three-Level Task Specification
This part introduces TLSQL’s three-level specification logic, as illustrated in Figure 4, where certain statements are optional. The workflow supports three levels of task specification with increasing degrees of control:
@@ 4. Conclusion
In this work, we presented TLSQL, a declarative SQL-like interface for table learning over modern relational databases. By translating TLSQL statements into standard SQL queries and structured task descriptions consumable by downstream table learning frameworks, TLSQL eliminates the need for data export, manual preprocessing, or low-level pipeline management. We believe TLSQL provides a compelling foundation for future research in database-centric learning, offering both a conceptual framework and a practical prototype. As future work, we plan to integrate large language models to generate and refine TLSQL programs from natural language, further simplifying end-to-end table learning in real-world settings.

### v3

@@ Abstract.
Table learning, which lies at the intersection of machine learning and modern database systems, has recently attracted growing attention. However, existing table learning frameworks typically require explicit data export and extensive feature engineering, creating a high barrier for database practitioners. We present TLSQL (Table Learning Structured Query Language), a system that enables table learning directly over relational databases via SQL-like declarative specifications. TLSQL is implemented as a lightweight Python library that translates these specifications into standard SQL queries and structured learning task descriptions. The generated SQL queries are executed natively by the database engine, while the task descriptions are consumed by downstream table learning frameworks. This design allows users to focus on modeling and analysis rather than low-level data preparation and pipeline orchestration. Experiments on real-world datasets demonstrate that TLSQL effectively lowers the barrier to integrating machine learning into database-centric workflows. Our code is available at https://github.com/rllm-project/tlsql/.
@@ 1. Introduction
In sum, by bridging relational databases and machine learning through a SQL-like interface, TLSQL lowers the barrier for database users to integrate machine learning into SQL-centric workflows and provides a practical foundation for database-centric learning systems.
@@ 3.3. Three-Level Task Specification
Figure 4 illustrates the three-level task specification mechanism of TLSQL. The convert_workflow_queries function accepts multiple TLSQL queries that jointly define a learning workflow. By analyzing the logical relationships among these queries, TLSQL automatically optimizes and consolidates them into a coherent execution plan. This design supports three levels of task specification with increasing degrees of user control:
@@ 4. Conclusion
In this work, we presented TLSQL, a declarative SQL-like interface for table learning over modern relational databases. TLSQL allows users to specify predictive tasks using simple constructs, PREDICT VALUE, TRAIN WITH, and VALIDATE WITH, eliminating the need for complex manual data export, feature engineering, or low-level pipeline management. As such, TLSQL enables database users to seamlessly integrate machine learning into existing SQL-centric workflows, significantly facilitating its deployment in real-world production environments. We believe TLSQL provides a strong foundation for future research in database-centric learning, offering both a conceptual framework and a practical prototype.